# Security of Language Models for Code: A Systematic Literature Review

### v1

@@ Abstract.
Language models for code (CodeLMs) have emerged as powerful tools for code-related tasks, outperforming traditional methods and standard machine learning approaches. However, these models are susceptible to security vulnerabilities, drawing increasing research attention from domains such as software engineering, artificial intelligence, and cybersecurity. Despite the growing body of research focused on the security of CodeLMs, a comprehensive survey in this area remains absent. To address this gap, we systematically review 67 relevant papers, organizing them based on attack and defense strategies. Furthermore, we provide an overview of commonly used language models, datasets, and evaluation metrics, and highlight open-source tools and promising directions for future research in securing CodeLMs.
@@
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.1122456ccs: Software and its engineering Automatic programmingccs: Security and privacy Software security engineering
@@ 1. Introduction
Language models for code (in short as CodeLMs), such as Codex (OpenAI, 2023) and CodeLlama (Meta, 2023), have become powerful tools in Software Engineering (SE) (Zhang et al., 2023a; Liu et al., 2024), addressing tasks such as code search (Sun et al., 2022, 2024d), code completion/generation (Svyatkovskiy et al., 2019; Izadi et al., 2024; Wang et al., 2024b), and code summarization (Sun et al., 2023b, 2024c). These language models utilize advanced deep learning techniques to analyze and represent code, significantly improving performance compared to traditional methods. One notable development is GitHub Copilot (GitHub, 2023), an AI-powered coding assistant built on the Codex (OpenAI, 2023) model. Trained on diverse programming patterns and vast open-source code from GitHub, Codex enables Copilot to suggest and complete code in real-time, greatly enhancing developers’ productivity.
@@
Despite the success of CodeLMs in solving various code-related tasks, their inherent vulnerabilities raise significant concerns, particularly in security-critical applications such as network security and malware detection. These models are susceptible to backdoor attacks, where hidden triggers are embedded during training (Sun et al., 2023a). Once activated by an attacker, the model’s behavior changes maliciously. For example, Schuster et al. (Schuster et al., 2021) demonstrated that backdoors in CodeLMs can significantly increase the likelihood of generating vulnerable code. Additionally, adversarial attacks (Szegedy et al., 2014; Chakraborty et al., 2021), which occur during the inference phase, add subtle perturbations to inputs, causing models to make incorrect predictions with high confidence, further highlighting the fragility of these models.
@@
To counter attacks on CodeLMs, various defense methods have been developed. For backdoor attacks, defenses often focus on detecting outliers in the training data (Ramakrishnan and Albarghouthi, 2022; Li et al., 2024b). After identifying and removing suspicious data points, the model is retrained on the cleaned dataset to ensure security. For adversarial attacks, adversarial training (Henkel et al., 2022) is an effective defense, where adversarial samples are incorporated into the training data to enhance the model’s robustness against such attacks.
@@
While numerous methods for attacking and defending CodeLMs have been proposed, a comprehensive review of the existing research in this field remains absent. To address this gap, we present a systematic survey on the security of CodeLMs. We conduct extensive literature searches across six widely-used databases using 407 search keywords, complemented by a snowballing approach to ensure thoroughness. This process yields 67 relevant papers. Based on these studies, we categorize common attacks on CodeLMs, focusing on backdoor and adversarial attacks, and summarize corresponding defense strategies. Additionally, we review common experimental settings, including datasets, language models, evaluation metrics, and artifact accessibility. Finally, we highlight open challenges and future opportunities in securing CodeLMs.
@@
  • First Comprehensive Survey on Security of CodeLMs. To the best of our knowledge, we conduct the first systematic literature review on the security of CodeLMs, analyzing 67 relevant papers. Our survey categorizes various attack and defense strategies for CodeLMs, covering multiple critical aspects and application scenarios.
@@
against CodeLMs] [ §5.2 Backdoor Defenses on CodeLMs [ §5.2.1 Overview of Backdoor Defenses] [ §5.2.2 Pre-training Backdoor Defenses] [ §5.2.3 In-training Backdoor Defenses] [ §5.2.4 Post-training Backdoor Defenses] [ §5.2.5 Summary of Backdoor Defenses] ] [ §5.3 Adversarial Defenses on CodeLMs [ §5.3.1 Overview of Adversarial Defenses] [ §5.3.2 Adversarial Training] [ §5.3.3 Model Modification] [ §5.3.4 Additional Models] [ §5.3.5 Summary of Adversarial Defenses] ] ] [ §6 ANSWER TO RQ3: EMPIRICAL
@@
SETTING AND EVALUATION [ §7.1 Experimental Datasets] [ §7.2 Experimental CodeLMs] [ §7.3 Evaluation Metrics] [ §7.4 Artifact Accessibility] ] [ §8 CHALLENGES AND OPPORTUNITIES] [ §9 CONCLUSION] ]
@@
Structure of the paper: The structure of this paper is presented in Figure 1. The subsequent sections of this paper are structured as follows. Section 2 introduces the background of language models for code and language model security. Section 3 presents the survey methodology that we adhere to. In After that, we present an overview of attacks and defenses against CodeLMs. Sections 4.2 and 4.3 provide a review of existing backdoor threats, adversarial threats in CodeLMs, respectively. On the other hand, Sections 5.2 and 5.3 provide a review of existing backdoor defenses and adversarial defenses. Section 6 reviews the existing empirical studies that explore the capabilities of attack or defense techniques and their impact on CodeLMs. Section 7 summarises the datasets, LMs, and metrics widely utilized in the field of CodeLM security. In Section 8, we discuss the challenges and opportunities in this field for future work. Finally, Section 9 provides a conclusion of this survey.
@@ 2. Background and Related Work
In this section, we introduce the background of CodeLMs, discuss security threats to neural models, and provide a comparison with existing related surveys.
@@ 2.1. Language Models for Code
CodeLMs refer to models that interpret code using neural language models (Palacio et al., 2023). In 2012, Hindle et al. pioneered the introduction of n-gram models into the SE field (Hindle et al., 2016). Since then, CodeLMs have illustrated performance in various code-related tasks such as code search (Gu et al., 2018; Sachdev et al., 2018; Haldar et al., 2020; Sun et al., 2022), code summarization (LeClair et al., 2019; Jiang et al., 2017; Haque et al., 2020; Ahmad et al., 2020; Sun et al., 2023c), and code completion (Alon et al., 2020; Karampatsis and Sutton, 2019; Kim et al., 2021; Svyatkovskiy et al., 2020). These studies shape the problem at hand as a text-to-text transformation, where the input and output of the model are both text strings (Mastropaolo et al., 2021). Furthermore, considering that code embodies rich structural information, some studies propose extracting the structural information of code as input to capture deeper features of the code (Mou et al., 2016; Zhang et al., 2019; Wan et al., 2019; Zeng et al., 2023). Later, with significant advancements in transfer learning in the field of NLP (e.g., GPT (Radford et al., 2018), BERT (Devlin et al., 2019), T5 (Raffel et al., 2020)), the fundamental idea involves initially pre-training a model on a large and generic dataset using a self-supervised task (e.g., masking tokens in strings and requiring the model to guess the masked tokens). Subsequently, the pre-trained model is fine-tuned on smaller and specialized datasets, each designed to support specific tasks (Han et al., 2021). In this context, Feng et al. (Feng et al., 2020) propose CodeBERT, which integrates a token prediction scheme to understand code by predicting subsequent tokens, thereby enhancing its understanding of programming languages for tasks such as code completion and error detection. Building upon this, Guo et al. (Guo et al., 2021) introduce GraphCodeBERT, which incorporates edge-type prediction to identify relationships between code elements and represent them as graphs. This enables GraphCodeBERT to leverage code structure, improving its performance in tasks like code summarization and program analysis. In recent years, researchers have observed that significantly increasing the scale of pre-trained models can enhance their capacity, leading to significant performance improvements when the parameter size surpasses a certain threshold (Hoffmann et al., 2022; Shanahan, 2024; Taylor et al., 2022). The term LLMs has been introduced to distinguish language models based on their parameter size, particularly referring to large program language models (Zhao et al., 2023; Zhang et al., 2024b). Specialized LLMs for code have emerged successively, such as GitHub Copilot (GitHub, 2023), Amazon’s CodeWhisperer (Amazon, 2023), the OpenAI Code Interpreter (OpenAI, 2023) integrated into ChatGPT, and Code Llama (Meta, 2023). This marks a new stage in CodeLMs. These LLMs for code trained on code-specific datasets represent not only incremental improvements but also a paradigm shift in code understanding, generation, and efficiency. They have laid the foundation for breakthroughs in performance across various code-related tasks.
@@ 2.2. Language Model Security
In recent years, DL models have been widely applied to real-world tasks such as text analysis (Chen et al., 2019c, a), image classification (Krizhevsky et al., 2012; He et al., 2016; Han et al., 2024a), speech recognition (Dahl et al., 2011; Hinton et al., 2012), and autonomous driving (Li et al., 2024a; Chen et al., 2019d; Guo et al., 2024). Despite the impressive performance of DL models, their security is of high concern. Due to the inherent flaws in DL models, attackers can craft a sample carefully to mislead DL models or guide the learner to train a poor model (Chakraborty et al., 2021).
For example, attackers can significantly degrade overall performance by tampering with sensor-manipulated training data, leading to targeted misclassifications or the insertion of backdoors (Zhu et al., 2019; Gu et al., 2017; Liu et al., 2018). He et al. (He et al., 2022) categorize the security threats to DL models into four types: model extraction attack threats, model inversion attack threats, data poisoning attack threats, and adversarial attack threats. Model extraction attacks attempt to replicate a DL model via the provided API without prior knowledge of its training data and algorithms (Tramèr et al., 2016). Such attacks not only compromise the confidentiality of the model and harm its owner’s interests but also allow for the construction of an approximately equivalent white-box model, facilitating further attacks (e.g., adversarial attacks) (Papernot et al., 2017). Model inversion attacks exploit the information flow extracted and abstracted from the training data by the model, using the model’s predictions or confidence scores to restore data memberships or data properties (Song et al., 2017). Poisoning attacks primarily aim to reduce the prediction accuracy of DL models by contaminating the training data, thus compromising the model’s availability (Jagielski et al., 2018; Han et al., 2024b). Recently, many researchers have leveraged poisoning attacks to create backdoors in the target model, known as backdoor attacks. While the model may behave normally most of the time, it will produce incorrect predictions when exposed to carefully crafted data. With the pre-implanted backdoor and trigger data, attackers can manipulate prediction results and initiate further attacks. Adversarial attacks add imperceptible perturbations to normal samples during the prediction process of DL models, generating adversarial examples (Carlini and Wagner, 2017). Adversarial examples must both deceive the classifier and be imperceptible to humans. Similar to poisoning attacks, adversarial attacks also cause the model to misclassify malicious samples. The difference is that poisoning attacks insert malicious samples into the training data, directly contaminating the model, while adversarial attacks exploit the model’s weaknesses using adversarial examples to achieve incorrect prediction results.
@@ 2.3. Comparison with Existing Work
Existing work has discussed various aspects of current CodeLMs, including their capabilities on different specific tasks (Lin et al., 2020; Ahmed et al., 2023; Zhang et al., 2024a), modeling source code with CodeLMs (Le et al., 2021), the non-functional properties of CodeLMs (Yang et al., 2024b; Wang et al., 2024a), the trojan security in CodeLMs (Hussain et al., 2023b), and so on. Among them, the works of (Yang et al., 2024b) and (Hussain et al., 2023b) also focus on the security aspects of CodeLMs, which is similar to our literature review. Specifically, Yang et al. (Yang et al., 2024b) explore seven non-functional properties of large CodeLMs (CodeLLMs, for short), including robustness, security, privacy, explainability, efficiency, and usability. For the security property, they discuss the potential security threats that CodeLLMs may face (i.e., CodeLLMs as the victim), primarily focusing on backdoor security threats. Hussain et al. (Hussain et al., 2023b) establish a taxonomy of trojan AI in code and its triggers based on research in explainable AI and trojan AI. Additionally, they drew actionable insights from explainable AI that are applicable to the field of trojan AI in code. However, the aforementioned studies focus on specific aspects of CodeLM security and do not provide a comprehensive overview of the current state of security research in CodeLMs. Overall, there is still a lack of a systematic review of CodeLM security. In this paper, we provide a systematic literature review on the security of CodeLMs, including backdoor and adversarial security, with the aim of providing a foundation for future research in this area.

### v2

@@ Abstract.
Language models for code (CodeLMs) have emerged as powerful tools for code-related tasks, outperforming traditional methods and standard machine learning approaches. However, these models are susceptible to security vulnerabilities, drawing increasing research attention from domains such as software engineering, artificial intelligence, and cybersecurity. Despite the growing body of research focused on the security of CodeLMs, a comprehensive survey in this area remains absent. To address this gap, we systematically review 68 relevant papers, organizing them based on attack and defense strategies. Furthermore, we provide an overview of commonly used language models, datasets, and evaluation metrics, and highlight open-source tools and promising directions for future research in securing CodeLMs.
@@
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.1122456journal: tosemjournalvolume: 0journalnumber: 0article: 1publicationmonth: 0ccs: Software and its engineering Automatic programmingccs: Security and privacy Software security engineering
@@ 1. Introduction
Language models for code (in short as CodeLMs), such as Codex (OpenAI, 2023) and CodeLlama (Meta, 2023), have become powerful tools in Software Engineering (SE) (Liu et al., 2024), addressing tasks such as code search (Sun et al., 2022, 2024b) and code completion (Svyatkovskiy et al., 2019; Izadi et al., 2024). These language models utilize advanced deep learning techniques to analyze and represent code, significantly improving performance compared to traditional methods. One notable development is GitHub Copilot (GitHub, 2023), an AI-powered coding assistant built on the Codex (OpenAI, 2023) model. Trained on diverse programming patterns and vast open-source code from GitHub, Codex enables Copilot to suggest and complete code in real-time, greatly enhancing developers’ productivity.
@@
Despite the success of CodeLMs in solving various code-related tasks, their inherent vulnerabilities raise significant concerns, particularly in security-critical applications such as network security and malware detection. These models are susceptible to backdoor attacks, where hidden triggers are embedded during training (Sun et al., 2023). Triggers are specific input patterns designed by attackers, such as particular variable names or dead code snippets in the code, which are used to activate the model’s malicious behavior. For example, Schuster et al. (Schuster et al., 2021) demonstrated that backdoors in CodeLMs can significantly increase the likelihood of generating vulnerable code. In addition to backdoor attacks, adversarial attacks (Szegedy et al., 2014; Chakraborty et al., 2021) present another challenge to the security of CodeLMs. These attacks occur during the inference phase, where attackers add subtle perturbations to inputs, causing models to make incorrect predictions with high confidence, further highlighting the fragility of these models.
@@
To counter attacks on CodeLMs, various defense methods have been developed. For backdoor attacks, defenses often focus on detecting outliers in the training data (Ramakrishnan and Albarghouthi, 2022; Li et al., 2024b; Sun et al., 2025a, b). After identifying and removing suspicious data points, the model is retrained on the cleaned dataset to ensure security. For adversarial attacks, adversarial training (Henkel et al., 2022) is an effective defense, where adversarial samples are incorporated into the training process to enhance the model’s robustness against such attacks.
@@
While numerous methods for attacking and defending CodeLMs have been proposed, a comprehensive review of the existing research in this field remains absent. To address this gap, we present a systematic survey on the security of CodeLMs. We conduct extensive literature searches across six widely-used databases using 407 search keywords, complemented by a snowballing approach to ensure thoroughness. This process yields 68 relevant papers. Based on these studies, we categorize common attacks on CodeLMs, focusing on backdoor and adversarial attacks, and summarize corresponding defense strategies. Additionally, we review common experimental settings, including datasets, language models, evaluation metrics, and artifact accessibility. Finally, we highlight open challenges and future opportunities in securing CodeLMs.
@@
  • First Comprehensive Survey on Security of CodeLMs. To the best of our knowledge, we conduct the first systematic literature review on the security of CodeLMs, analyzing 68 relevant papers. Our survey categorizes various attack and defense strategies for CodeLMs, covering multiple critical aspects and application scenarios.
@@
against CodeLMs] [ §5.2 Backdoor Defenses on CodeLMs [ §5.2.1 Overview of Backdoor Defenses
against CodeLMs] [ §5.2.2 Pre-training Backdoor Defenses] [ §5.2.3 In-training Backdoor Defenses] [ §5.2.4 Post-training Backdoor Defenses] [ §5.2.5 Summary of Backdoor Defenses] ] [ §5.3 Adversarial Defenses on CodeLMs [ §5.3.1 Overview of Adversarial Defenses
against CodeLMs] [ §5.3.2 Input Modification] [ §5.3.3 Model Modification] [ §5.3.4 Additional Models] [ §5.3.5 Summary of Adversarial Defenses] ] ] [ §6 ANSWER TO RQ3: EMPIRICAL
@@
SETTING AND EVALUATION [ §7.1 Experimental Datasets] [ §7.2 Experimental CodeLMs] [ §7.3 Evaluation Metrics] [ §7.4 Artifact Accessibility] ][ §8 CHALLENGES AND OPPORTUNITIES [ §8.1 Challenges and Opportunities in
Attacks against CodeLMs] [ §8.2 Challenges and Opportunities in
Defenses against Attacks on CodeLMs] ] [ §9 THREATS TO VALIDITY] [ §10 CONCLUSION] ]
@@
Structure of the paper: The structure of this paper is presented in Figure 1. The subsequent sections of this paper are structured as follows. Section 2 introduces the background of language models for code and language model security. Section 3 presents the survey methodology that we adhere to. After that, we present an overview of attacks and defenses against CodeLMs. Sections 4.2 and 4.3 provide a review of existing backdoor threats, adversarial threats in CodeLMs, respectively. On the other hand, Sections 5.2 and 5.3 provide a review of existing backdoor defenses and adversarial defenses. Section 6 reviews the existing empirical studies that explore the capabilities of attack or defense techniques and their impact on CodeLMs. Section 7 summarises the datasets, LMs, and metrics widely utilized in the field of CodeLM security. In Section 8, we outline the challenges and opportunities in this field for future work. In Section 9, we discuss the potential threats to the validity of this study. Finally, Section 10 provides a conclusion of this survey.
@@ 2. Background and Related Work
Table 1. Common technical terms related to the security of CodeLMs.
Terms Explanation
CodeLMs Language models used for understanding, generating, or processing code, leveraging neural networks and deep learning technique similar to NLP. Examples include CodeBERT, CodeT5, Codex, and DeepSeek Coder.
Trigger A trigger is a specific input pattern carefully designed by an attacker. It can be a specific token, such as variable names, code snippets, or words in comments. The attacker typically embed these triggers into the model during training, ensuring that the model behaves normally under regular inputs but exhibits specific anomalous behaviors when encountering the trigger.
Benign Sample A benign sample (also known as a clean sample) is a data sample that has not been affected by malicious modifications. It is considered “normal” and does not contain harmful attributes such as backdoor triggers or adversarial perturbations.
Poisoned Sample A poisoned sample is a maliciously modified data sample that contains attacker-designed triggers, which are used to implant backdoors into the model during training.
Benign Model A Benign model (also known as a clean model) is a model trained on Benign samples, behaving as expected without backdoors or malicious features.
Poisoned Model A poisoned model is a model trained on poisoned samples, implanted with a backdoor by an attacker. It behaves normally under regular conditions but produces malicious outputs when exposed to a trigger.
Adversarial Perturbation An adversarial perturbation is a minor modification to input data, such as variable name changes or dead code insertion, designed to mislead the model into making incorrect predictions without significantly altering the data’s meaning.
Adversarial Example Input data altered by adding small perturbations that mislead the model into making incorrect predictions, typically imperceptible to humans but misleading to the model.
Source Label The source label (also known as the victim label) refers to the original label prediction by the model for an input before the activation of a backdoor attack or an adversarial attack.
Target Label The target label is the label that the attacker wants the model to output, aiming to misclassify input samples containing triggers or adversarial perturbations into the attacker’s specified category.
Untargeted Attack An untargeted attack aims to cause the model’s prediction to deviate from the original correct class without enforcing classification into a specific incorrect label.
Targeted Attack A targeted attack aims to manipulate the model’s prediction, causing the input sample to be misclassified into a specific class predetermined by the attacker.
In this section, we introduce the background of CodeLMs, discuss security threats to neural models, and provide a comparison with existing related surveys. To enhance readability and facilitate understanding, we first provide brief descriptions and explanations of common technical terms related to the security of CodeLMs, as shown in Table 1.
@@ 2.1. Language Models for Code
CodeLMs refer to models that interpret code using neural language models (Palacio et al., 2023). In 2012, Hindle et al. pioneered the introduction of n-gram models into the SE field (Hindle et al., 2016). Since then, CodeLMs have illustrated performance in various code-related tasks such as code search (Gu et al., 2018; Sachdev et al., 2018; Haldar et al., 2020; Sun et al., 2022), code summarization (LeClair et al., 2019; Jiang et al., 2017; Haque et al., 2020; Ahmad et al., 2020; Fang et al., 2024; Sun et al., 2024c), and code completion (Alon et al., 2020; Karampatsis and Sutton, 2019; Kim et al., 2021; Svyatkovskiy et al., 2020). These studies shape the problem at hand as a text-to-text transformation, where the input and output of the model are both text strings (Mastropaolo et al., 2021). Furthermore, considering that code embodies rich structural information, some studies propose extracting the structural information of code as input to capture deeper features of the code (Mou et al., 2016; Zhang et al., 2019; Wan et al., 2019; Zeng et al., 2023). Later, with significant advancements in transfer learning in the field of NLP (e.g., GPT (Radford et al., 2018), BERT (Devlin et al., 2019), T5 (Raffel et al., 2020)), the fundamental idea involves initially pre-training a model on a large and generic dataset using a self-supervised task (e.g., masking tokens in strings and requiring the model to guess the masked tokens). Subsequently, the pre-trained model is fine-tuned on smaller and specialized datasets, each designed to support specific tasks (Han et al., 2021). In this context, Feng et al. (Feng et al., 2020) propose CodeBERT, which integrates a token prediction scheme to understand code by predicting subsequent tokens, thereby enhancing its understanding of programming languages for tasks such as code completion and error detection. Building upon this, Guo et al. (Guo et al., 2021) introduce GraphCodeBERT, which incorporates edge-type prediction to identify relationships between code elements and represent them as graphs. This enables GraphCodeBERT to leverage code structure, improving its performance in tasks like code summarization and program analysis. In recent years, researchers have observed that significantly increasing the scale of pre-trained models can enhance their capacity, leading to significant performance improvements when the parameter size surpasses a certain threshold (Hoffmann et al., 2022; Shanahan, 2024; Taylor et al., 2022). The term LLMs has been introduced to distinguish language models based on their parameter size, particularly referring to large program language models (Zhao et al., 2023). Specialized LLMs for code have emerged successively, such as GitHub Copilot (GitHub, 2023), Amazon’s CodeWhisperer (Amazon, 2023), the OpenAI Code Interpreter (OpenAI, 2023) integrated into ChatGPT, and Code Llama (Meta, 2023). This marks a new stage in CodeLM. These LLMs for code trained on code-specific datasets represent not only incremental improvements but also a paradigm shift in code understanding, generation, and efficiency. They have laid the foundation for breakthroughs in performance across various code-related tasks. For example, Fried et al. (Fried et al., 2023) introduce a large language model, InCoder, and try zero-shot training on the CodeXGLUE (Lu et al., 2021) Python dataset and achieve impressive results. Wu et al. (Wu et al., 2024) propose the first LLM-driven efficient and effective framework for detecting Ponzi smart contracts written in the Solidity code.
@@ 2.2. Language Model Security
In recent years, DL models have been widely applied to real-world tasks such as text analysis (Chen et al., 2019c, a), image classification (Krizhevsky et al., 2012; He et al., 2016), speech recognition (Dahl et al., 2011; Hinton et al., 2012), and autonomous driving (Li et al., 2024a; Chen et al., 2019d). Despite the impressive performance of DL models, their security is of high concern. Due to the inherent flaws in DL models, attackers can craft a sample carefully to mislead DL models or guide the learner to train a poor model (Chakraborty et al., 2021). For example, attackers can significantly degrade overall performance by tampering with sensor-manipulated training data, leading to targeted misclassifications or the insertion of backdoors (Zhu et al., 2019; Gu et al., 2017; Liu et al., 2018). He et al. (He et al., 2022) categorize the security threats to DL models into four types: model extraction attack threats, model inversion attack threats, data poisoning attack threats, and adversarial attack threats. Model extraction attacks attempt to replicate a DL model via the provided API without prior knowledge of its training data and algorithms (Tramèr et al., 2016). Such attacks not only compromise the confidentiality of the model and harm its owner’s interests but also allow for the construction of an approximately equivalent white-box model, facilitating further attacks (e.g., adversarial attacks) (Papernot et al., 2017). Model inversion attacks exploit the information flow extracted and abstracted from the training data by the model, using the model’s predictions or confidence scores to restore data memberships or data properties (Song et al., 2017). Poisoning attacks primarily aim to reduce the prediction accuracy of DL models by contaminating the training data, thus compromising the model’s availability (Jagielski et al., 2018). For example, Lv et al. (Lv et al., 2023) propose a generic data-free backdoor poisoning attack against diverse DL tasks and models by utilizing a substitute dataset irrelevant to the main task. Recently, many researchers have leveraged poisoning attacks to create backdoors in the target model, known as backdoor attacks. While the model may behave normally most of the time, it will produce incorrect predictions when exposed to carefully crafted data. With the pre-implanted backdoor and trigger data, attackers can manipulate prediction results and initiate further attacks. Adversarial attacks add imperceptible perturbations to normal samples during the prediction process of DL models, generating adversarial examples (Carlini and Wagner, 2017). Adversarial examples must both deceive the classifier and be imperceptible to humans. Similar to poisoning attacks, adversarial attacks also cause the model to misclassify malicious samples. The difference is that poisoning attacks insert malicious samples into the training data, directly contaminating the model, while adversarial attacks exploit the model’s weaknesses using adversarial examples to achieve incorrect prediction results.
@@ 2.3. Comparison with Existing Work
Existing work has discussed various aspects of current CodeLMs, including their capabilities on different specific tasks (Lin et al., 2020; Ahmed et al., 2023; Zhang et al., 2024a), the impact of data on CodeLMs (Zhuo et al., 2023), modeling source code with CodeLMs (Le et al., 2021), the non-functional properties of CodeLMs (Yang et al., 2024b; Wang et al., 2024), the trojan security in CodeLMs (Hussain et al., 2023b), and so on. Among them, the works of (Yang et al., 2024b; Hussain et al., 2023b; Zhuo et al., 2023) also focus on the security aspects of CodeLMs, which is similar to the focus of our literature review. Specifically, Yang et al. (Yang et al., 2024b) explore seven non-functional properties of large CodeLMs (CodeLLMs, for short), including robustness, security, privacy, explainability, efficiency, and usability. For the security property, they discuss the potential security threats that CodeLLMs may face (i.e., CodeLLMs as the victim), primarily focusing on backdoor security threats. Hussain et al. (Hussain et al., 2023b) establish a taxonomy of trojan AI in code and its triggers based on research in explainable AI and trojan AI. Additionally, they drew actionable insights from explainable AI that are applicable to the field of trojan AI in code. Zhuo et al. (Zhuo et al., 2023) conduct a comprehensive survey on data augmentation methods specifically designed for source code. They introduce various data augmentation techniques, discuss common strategies for improving data quality, and review the application of data augmentation in common source code tasks. They also emphasize the significant role of data augmentation in different source code scenarios, particularly in generating adversarial examples to enhance the performance, robustness, and security of CodeLMs. However, the aforementioned studies focus on specific aspects of CodeLM security and do not provide a comprehensive overview of the current state of security research in CodeLMs. Overall, there is still a lack of a systematic review of CodeLM security. In this paper, we provide a systematic literature review on the security of CodeLMs, including backdoor and adversarial security, with the aim of providing a foundation for future research in this area.